adovbush:

Determine if the triangle is a right triangle: 15 ft, 8 ft, 5 ft

adovbush:

@Ultrilliam @Jaynator495

Eiwoh2:

@ThisGirlPretty @Elsa213 Some people that can help.

adovbush:

Thanks

random231:

whats 15 squared?

random231:

whats 8 squared and 5 squared?

adovbush:

Idk

random231:

well then, idk how to help you sir.

adovbush:

Then why even come here

random231:

why even study math when you dont know what 5 squared is?

adovbush:

Who said I study math?

random231:

ohh my bad. you just copy answers from here

random231:

anyways farewell

Shadow:

A right triangle will follow the following rule \[a^2 + b^2 = c^2\] Where 'a' and 'b' are the legs and 'c' is the hypotenuse  The hypotenuse is always the side across from the right angle, or the 90 degree angle  The length of the hypotenuse is always greater than the sides. This is because the sum of the square of both the sides is equal to the square of the hypotenuse. So we know that 5 and 8 are the "legs." 15 must be the hypotenuse. But we must test if it is mathematically correct. So, \[5^2 + 8^2 = 15^2\] \[25 + 64 = 225\] \[89 = 225\] This is not equal. Therefore these measurements are not that of a right triangle. Also if you don't know what "squaring" is, it's basically this: \[15^2 = 15 \times 15\] \[8^2 = 8 \times 8\] \[5^2 = 5 \times 5\]

adovbush:

Thanks a lot